# CauScien: Uncovering Causality in Science

### v1

@@ 1 Introduction
Causal discovery in event sequences is a central problem across various domains, including cybersecurity [21], healthcare [32, 12], flight operations [19], and vehicle defects [28]. These sequences of discrete events $x_{i}$ recorded asynchronously over time often lead to outcomes (e.g, a diagnosed defect, a disease) denoted as labels $\boldsymbol{y}$. While becoming more available at scale, they remain challenging to interpret beyond associations. Understanding why specific events lead to particular outcomes is vital for effective diagnosis, prediction, and overall decision making [18, 29].
However, the majority of existing causal discovery methods remain computationally intractable in high dimensions [9, 11] with thousands of different nodes. Additionally,
practitioners frequently reason about causality within individual unknown sequences. For instance, ”what series of events captured by diagnostics led to this vehicle failure”.
@@
In this work, we introduce OSCAR: the first One-Shot multi-label Causal AutoRegressive discovery method. It leverages two Transformers [39] as density estimators to extract a compact interpretable subgraph with quantified causal indicators between events and labels, providing better explainability. Unlike traditional causal discovery methods that suffer label cardinality-dependent time complexity [17, 44, 11, 9], OSCAR supports causal discovery across thousands of nodes. Thanks to its fully parallelised structure, it provides sequence-specific explainability in a matter of minutes. We validate our approach on a real-world vehicular dataset comprising 29,100 event types as diagnosis trouble codes and 474 labels as error patterns (EPs) representing vehicle defects [22].
@@ 2 Related Work
Event sequences, such as diagnostic trouble codes in vehicles [28, 22] or electronic health records [32, 16], are often represented as a series of time-stamped discrete events $S=\{(t_{1},x_{1}),\ldots,(t_{L},x_{L})\}$ where $0\leq t_{1}<\ldots\leq t_{L}$ the time of occurrence of event type $x_{i}\in\mathbb{X}$ drawn from a finite vocabulary $\mathbb{X}$. In multi-label settings, a binary label vector $\boldsymbol{y}\in\{0,1\}^{|\mathbb{Y}|}$ is attached to $S$ and denotes the presence of multiple outcome labels drawn from $\mathbb{Y}$ occuring at final time step $t_{L}$. Forming a multi-labeled sequence $S_{l}=(S,(\boldsymbol{y}_{L},t_{L}))$.
Transformers have recently shown strong performance in high-dimensional event spaces for next-event or label prediction [39, 31, 36], including dual-architecture setups predicting both events and outcomes [22]—a structure we repurpose for causal discovery.
Neural autoregressive density estimators (NADEs) [1] factorise sequence likelihood via the chain rule, and modern Transformer-based NADEs have been applied to causal inference [8, 14] by simulating interventions or approximating Bayesian networks [8, 14].
Transformers as causal learners have been explored in sequential settings, with attention patterns interpreted as latent causal graphs [24, 23]. We extend the one-shot sequence-to-graph idea [33] to multi-label causal discovery to scale to tens of thousands of event types.
Multi-label Causal Discovery seeks to identify the Markov Boundary (MB) of each label—its minimal set of parents, children, and spouses—such that the label is conditionally independent of all other variables given its MB [37]. While classical constraint-based algorithms have shown success on low-dimensional tabular data [34, 44], their application to event sequences with multi-label outputs remains challenging due to dimensionality, sparsity, temporal dependencies, and distributional assumptions [9, 2, 44].
@@
Working with causal structure learning from observed data requires several assumptions, notably the causal Markov assumptions [26] states that a variable is conditionally independent of its non-descendants given its parents. We assume the following: an event is allowed to influence any future events only (temporal precedence A1), event lagged effects are contained in a fixed windows (A2), the transformers model perfectly the joint probability distribution of event and labels (A4) and causal sufficiency (A3). A discussion on the impact of assumptions is provided in Appendix H.
@@ 3 Methodology
We model each multi-labeled event sequence $S_{l}=(S,\boldsymbol{y}_{L})$ as a sequential Bayesian Network (Def. 1), over events $X_{i}\in\mathbb{X}$ and labels $Y_{j}\in\mathbb{Y}$ (Fig. 4). Our goal is to recover the MB of each $Y_{j}$. Specifically, we would like to assess how much additional information event $X_{i}$ occurring at step $i$ provides about label $Y_{j}$ when we already know the past sequence of events $\boldsymbol{Z}=S_{<i}$. We essentially try to answer if:
@@
where $D_{KL}$ denotes the Kullback-Leibler divergence [3]. The distributional difference between the conditionals $P(Y_{j}|X_{i},\boldsymbol{Z}),P(Y_{j}|\boldsymbol{Z})$ is akin to Information Gain $I_{G}$ [30] conditioned on past events:
$I_{G}(x_{i},Y_{j}|z_{i})\triangleq D_{KL}(P(Y_{j}|X_{i}=x_{i},\boldsymbol{Z}=z_{i}))||P(Y_{j}|Z=z_{i}))$ (1)
Which is equals to the difference between the conditional entropies [3, 30] denoted as $H$:
$I_{G}(Y_{j},x_{i}|z_{i})=H(Y_{j}|z_{i})-H(Y_{j}|x_{i},z_{i})$ (2)
More generally, we can use the conditional mutual information (CMI) [3] as $I$ to assess conditional independence (Def. 4). It is simply the expected value of the information gain $I_{G}(Y_{j},x_{i}|z_{i})$ such as:
$I(Y_{j},X_{i}|\boldsymbol{Z})\triangleq H(Y_{j}|\boldsymbol{Z})-H(Y_{j}|\boldsymbol{Z},X_{i})=\mathbb{E}_{x_{i},z_{i}}[I_{G}(Y_{j},X_{i}=x_{i}|\boldsymbol{Z}=z_{i})])$ (3)
@@
Which can be interpreted as the expected value over all possible sequences $\boldsymbol{Z}$ of the deviation from independence of $X_{i},Y_{j}$ in this context.
In practice a per-label threshold $\theta_{j}\approx 0$ is applied to Eq (3) to identify conditional independence:
$Y_{j}\not\!\perp X_{i}\mid\boldsymbol{Z}\quad\Leftrightarrow\quad I(X_{i},Y_{j}\mid\boldsymbol{Z})>\theta_{j}\approx 0$ (4)
$\theta_{j}$ is dynamically computed for each label based on the mean and standard deviation of the CMI values across the sequence such that: $\theta_{j}=\mu_{Y_{j}}+k\cdot\sigma_{Y_{j}}$. We analyse the effect of $k$ in Fig. 3 as well as the effect of the sample type and number of sample contexts in Appendix G.2 and G.3.
We reuse the two architectures introduced by Math et al. [22] to perform next event prediction (CarFormer as $\text{Tf}_{x}$) and next labels (EPredictor as $\text{Tf}_{y}$) with past events $\boldsymbol{Z}=(x_{1},\cdots,x_{i-1})$
@@
Intuitively, Theorem 1 states that if our Transformers perfectly approximate the true joint distribution, then testing conditional mutual information at each step is sufficient to recover the Markov Boundary of each label sequentially. By induction, we prove that with bounded lagged effects of the previous events, we can restrict their causal influence and recover the correct MB of each label.
These guarantees rest on strong assumptions—causal sufficiency, oracle-quality density estimators, and bounded lagged effects. While these are unlikely to strictly hold in practice, they simplify identifiability and highlight where practical approximations may degrade performance (Appendix H)
A key advantage of our approach is its scalability. Unlike traditional methods whose complexity depends on the event and label cardinality $|\mathbb{X}|$ and $|\mathbb{Y}|$ [17], our method is agnostic to both. Fig. 1 shows all parallelised steps on GPUs. CMI estimations are independently performed for all positions $i\in[c,L]$, with the sampling pushed into the batch dimension and results averaged across labels. This transitions the time complexity from $\mathcal{O}(\text{BS}\times N\times L)$ to $\mathcal{O}(1)$ per batch. With $L$, the sequence length.
While deterministic DAGs reveal structural dependencies, they often obscure the magnitude and direction of influence between variables. Given that we can estimate conditional distributions, we define the causal indicator $\mathcal{C}\in[-1,1]$ [6, 4] between an event $X_{i}$ and a label $Y_{j}$ under context $\boldsymbol{Z}$ that we assume fixed for every measurement [6] as:
$\mathcal{C}(Y_{j},X_{i};\boldsymbol{Z}):=\mathbb{E}_{Z}[P(Y_{j}\mid X_{i},Z)-P(Y_{j}\mid Z)]$ (7)
This enables easy interpretation, for instance, if $\mathcal{C}<0$, then $X_{i}$ inhibits the occurrence of $Y_{j}$. We employ the term causal indicator to separate from causal strength measures, which, if using this formulation, can be problematic [15]. Ours serves more as an indication of the rise in label likelihood after observing a certain event, rather than a strength which is here the CMI.
@@ 4 Empirical Evaluation
Comparisons. Although no existing method directly targets one-shot multi-label causal discovery [9], we benchmark OSCAR against local structure learning (LSL) algorithms that estimate global Markov Boundaries. This includes established approaches such as CMB [7], MB-by-MB [40], PCD-by-PCD [43], IAMB [38] from the PyCausalFS package [44], as well as the more recent, state-of-the-art MI-MCF [20]. We used a $g4dn.12xlarge$ instance from AWS Sagemaker to run comparisons, containing 4 T4 GPUs. We used a combination of F1-Score, Precision, and Recall with different averaging [45] (Appendix  F.1) to perform the comparisons. The code for OSCAR, $\text{Tf}_{x},\text{Tf}_{y}$ and the evaluation are provided anonymously  [^1] https://github.com/Mathugo/OSCAR-One-Shot-Causal-AutoRegressive-discovery.git as well as the anonymised version of the dataset for reproducibility purposes. An Ablation of the NADEs quality is given in G.1
Vehicle Event Sequences Dataset. We evaluated our method on a real-world vehicular test set of $n=300,000$ sequences. It contains $|\mathbb{Y}|=474$ different error patterns and about $|\mathbb{X}|=29,100$ different DTCs forming sequences of $\approx 150\pm 90$ events. We used 105m backbones as $\text{Tf}_{x},\text{Tf}_{y}$ [22]. The two NADEs were not exposed to the test set during. The error patterns are manually defined by domain experts as boolean rules between DTCs in Eq (8) where $(y_{1})$ is a boolean definition based on diagnosis trouble codes $(x_{i})$:
@@
Table 1 shows comparision with $n=50,000$. We found out LSL algorithms failed to compute the Markov Boundaries within multiple days (3-day timeout), far exceeding practical limits for deployment. OSCAR, on the other hand, shows robust classification over a large amount of events $(29,100)$, especially $55\%$ precision, in a matter of minutes. This behaviour highlights the current infeasibility of multi-label causal discovery in high-dimensional event sequences, since it relies on expensive global CI-testings [9]. This positions OSCAR as a more feasible approach for large-scale causal per-sequence causal reasoning in production environments.

### v2

@@ 1 Introduction
Causal discovery in event sequences is a central problem across various domains, including cybersecurity [MANOCCHIO2024122564], healthcare [MedBERT, bihealth], flight operations [flight_service_cd], and vehicle defects [pdm_dtc_feature_extraction]. These sequences of discrete events $x_{i}$ recorded asynchronously over time often lead to outcomes (e.g, a diagnosed defect, a disease) denoted as labels $\boldsymbol{y}$. While becoming more available at scale, they remain challenging to interpret beyond associations. Understanding why specific events lead to particular outcomes is vital for effective diagnosis, prediction, and overall decision making [liu2025learning, shtp].
However, the majority of existing causal discovery methods remain computationally intractable in high dimensions [cd_temporaldata_review, hasan2023a] with thousands of different nodes. Additionally, practitioners frequently reason about causality within individual unknown sequences. For instance, ”what series of events captured by diagnostics led to this vehicle failure”.
@@
In this work, we introduce OSCAR: the first One-Shot multi-label Causal AutoRegressive discovery method. It leverages two Transformers [tf] as density estimators to extract a compact interpretable subgraph with quantified causal indicators between events and labels, providing better explainability. Unlike traditional causal discovery methods that suffer label cardinality-dependent time complexity [feature_selection_review, causality_based_feature_selection_2019, hasan2023a, cd_temporaldata_review], OSCAR supports causal discovery across thousands of nodes. Thanks to its fully parallelised structure, it provides sequence-specific explainability in a matter of minutes. We validate our approach on a real-world vehicular dataset comprising 29,100 event types as diagnosis trouble codes and 474 labels as error patterns (EPs) representing vehicle defects [math2024harnessingeventsensorydata].
@@ 2 Related Work
Event sequences, such as diagnostic trouble codes in vehicles [pdm_dtc_feature_extraction, math2024harnessingeventsensorydata] or electronic health records [MedBERT, pmlr-v219-labach23a], are often represented as a series of time-stamped discrete events $S=\{(t_{1},x_{1}),\ldots,(t_{L},x_{L})\}$ where $0\leq t_{1}<\ldots\leq t_{L}$ the time of occurrence of event type $x_{i}\in\mathbb{X}$ drawn from a finite vocabulary $\mathbb{X}$. In multi-label settings, a binary label vector $\boldsymbol{y}\in\{0,1\}^{|\mathbb{Y}|}$ is attached to $S$ and denotes the presence of multiple outcome labels drawn from $\mathbb{Y}$ occuring at final time step $t_{L}$. Forming a multi-labeled sequence $S_{l}=(S,(\boldsymbol{y}_{L},t_{L}))$.
Transformers have recently shown strong performance in high-dimensional event spaces for next-event or label prediction [tf, gpt, touvron2023llamaopenefficientfoundation], including dual-architecture setups predicting both events and outcomes [math2024harnessingeventsensorydata]—a structure we repurpose for causal discovery.
Neural autoregressive density estimators (NADEs) [NIPS1999_e6384711] factorise sequence likelihood via the chain rule, and modern Transformer-based NADEs have been applied to causal inference [density_estimator_2O21_journal_ci, im2024usingdeepautoregressivemodels] by simulating interventions or approximating Bayesian networks [density_estimator_2O21_journal_ci, im2024usingdeepautoregressivemodels].
Transformers as causal learners have been explored in sequential settings, with attention patterns interpreted as latent causal graphs [tf_learns_gradient_structure, make1010019]. We extend the one-shot sequence-to-graph idea [tf_causalinterpretation_neurips_2023] to multi-label causal discovery to scale to tens of thousands of event types.
Multi-label Causal Discovery seeks to identify the Markov Boundary (MB) of each label—its minimal set of parents, children, and spouses—such that the label is conditionally independent of all other variables given its MB [optimal_feature_set_cd]. While classical constraint-based algorithms have shown success on low-dimensional tabular data [constrainct_based_cd, causality_based_feature_selection_2019], their application to event sequences with multi-label outputs remains challenging due to dimensionality, sparsity, temporal dependencies, and distributional assumptions [cd_temporaldata_review, bn_np_hard, causality_based_feature_selection_2019].
@@
Working with causal structure learning from observed data requires several assumptions, notably the causal Markov assumptions [pearl_1998_bn] states that a variable is conditionally independent of its non-descendants given its parents. We assume the following: an event is allowed to influence any future events only (temporal precedence A1), event lagged effects are contained in a fixed windows (A2), the transformers model perfectly the joint probability distribution of event and labels (A4) and causal sufficiency (A3). A discussion on the impact of assumptions is provided in Appendix H.
@@ 3 Methodology
Conditional Mutual Information Estimation via Autoregressive Models. We model each multi-labeled event sequence $S_{l}=(S,\boldsymbol{y}_{L})$ as a sequential Bayesian Network (Def. 1), over events $X_{i}\in\mathbb{X}$ and labels $Y_{j}\in\mathbb{Y}$ (Fig. 5). Our goal is to recover the MB of each $Y_{j}$. Specifically, we would like to assess how much additional information event $X_{i}$ occurring at step $i$ provides about label $Y_{j}$ when we already know the past sequence of events $\boldsymbol{Z}=S_{<i}$. We essentially try to answer if:
@@
where $D_{KL}$ denotes the Kullback-Leibler divergence [cover1999elements]. The distributional difference between the conditionals $P(Y_{j}|X_{i},\boldsymbol{Z}),P(Y_{j}|\boldsymbol{Z})$ is akin to Information Gain $I_{G}$ [quinlan:induction] conditioned on past events:
$I_{G}(Y_{j},x_{i}|z)\triangleq D_{KL}(P(Y_{j}|X_{i}=x_{i},\boldsymbol{Z}=z))||P(Y_{j}|\boldsymbol{Z}=z))$ (1)
Which is equals to the difference between the conditional entropies [cover1999elements, quinlan:induction] denoted as $H$:
$I_{G}(Y_{j},x_{i}|z)=H(Y_{j}|z)-H(Y_{j}|x_{i},z)$ (2)
More generally, we can use the conditional mutual information (CMI) [cover1999elements] as $I$ to assess conditional independence (Def. 4). It is simply the expected value of the information gain $I_{G}(Y_{j},x_{i}|z)$ such as:
$I(Y_{j},X_{i}|\boldsymbol{Z})\triangleq H(Y_{j}|\boldsymbol{Z})-H(Y_{j}|\boldsymbol{Z},X_{i})=\mathbb{E}_{p(z)}[I_{G}(Y_{j},X_{i}|\boldsymbol{Z}=z)]$ (3)
@@
It can be interpreted as the expected value over all possible contexts $\boldsymbol{Z}$ of the deviation from independence of $X_{i},Y_{j}$ in this context. To approximate Eq. (3), a naive Monte Carlo [Doucet2001] approximation is performed where we draw $N$ random variations of the conditioning set $z^{(l)}=\{x^{(l)}_{0},\ldots,x^{(l)}_{i-1}\}$, denoting the $l$-th sampled particle:
$\hat{I}_{N}(Y_{j},X_{i}\mid\boldsymbol{Z})$ $=\frac{1}{N}\sum_{l=1}^{N}I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z^{(l)})$ (4)
This estimator is unbiased because the contexts $z^{(l)}$ are sampled directly from $\text{Tf}_{x}$ using a proposal $Q$ with the same support as $P(\boldsymbol{Z})$. Since $I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z)$ is a difference between conditional entropies ((2)), it is thus bounded uniformly [cover1999elements] by the log of supports such as:
$0<I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z^{(l)})=H(Y_{j}|z^{(l)})-H(Y_{j}|X_{i},z^{(l)}))\leq H(Y_{j})\leq\log{|\mathbb{Y}|}$
Thus the posterior variance of $f_{i}=I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z^{(l)})$ satisfies $\sigma^{2}_{f_{i}}\triangleq\mathbb{E}_{p(z)}[f^{2}_{i}(p(z)]-I^{2}(f_{i})<+\infty$ [Doucet2001] then the variance of $\hat{I}_{N}(f_{i})$) is equal to $\textit{var}(\hat{I}_{N}(f_{i}))=\frac{\sigma^{2}_{f_{t}}}{N}$ and from the strong law of large numbers:
$\hat{I}_{N}$ $\xrightarrow[N\to+\infty]{\text{a.s.}}\mathbb{E}_{p(z)}\!\left[I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z)\right]\triangleq I(f_{i}).$ (5)
Sequential Markov Boundary Recovery. In practice a per-label threshold $\theta_{j}\approx 0$ is applied to Eq (4) to identify conditional independence:
$Y_{j}\not\!\perp X_{i}\mid\boldsymbol{Z}\quad\Leftrightarrow\quad I(Y_{j},X_{i}\mid\boldsymbol{Z})>\theta_{j}\approx 0$ (6)
$\theta_{j}$ is dynamically computed for each label based on the mean and standard deviation of the CMI values across the sequence such that: $\theta_{j}=\mu_{Y_{j}}+k\cdot\sigma_{Y_{j}}$. We analyse the effect of $k$ in Fig. 4 as well as the effect of the proposal $Q$ and number of particles in Appendix G.2 and G.3.
We reuse the two architectures introduced by math2024harnessingeventsensorydata to perform next event prediction (CarFormer as $\text{Tf}_{x}$) and next labels (EPredictor as $\text{Tf}_{y}$) with past events $\boldsymbol{Z}=(x_{1},\cdots,x_{i-1})$
@@
Intuitively, Theorem 1 states that if our Transformers perfectly approximate the true joint distribution, then testing conditional mutual information at each step is sufficient to recover the Markov Boundary of each label sequentially. By induction, we prove that with bounded lagged effects of the previous events, we can restrict their causal influence and recover the correct MB of each label in the associated Bayesian Network (Fig. 5).
These guarantees rest on strong assumptions—causal sufficiency, oracle-quality density estimators, and bounded lagged effects. While these are unlikely to strictly hold in practice, they simplify identifiability and highlight where practical approximations may degrade performance (Appendix H). We particularly observed that for rare labels, the one-shot performances drastically dropped. We analyze this behavior in the experiment section in Fig. 2. One must carefully evaluate $\text{Tf}_{y}$ classification performance on downstream tasks using macro metrics before applying OSCAR.
Computation. A key advantage of our approach is its scalability. Unlike traditional methods whose complexity depends on the event and label cardinality $|\mathbb{X}|$ and $|\mathbb{Y}|$ [feature_selection_review], our method is agnostic to both. Fig. 1 shows all parallelised steps on GPUs. CMI estimations are independently performed for all positions $i\in[c,L]$, with the sampling pushed into the batch dimension and results averaged across particles using Eq. 4. This transitions the time complexity from $\mathcal{O}(\text{BS}\times N\times L)$ to $\mathcal{O}(1)$ per batch. A Pytorch [pytorch] implementation is given in Appendix K.
Causal Indicator. While deterministic DAGs reveal structural dependencies, they often obscure the magnitude and direction of influence between variables. Given that we can estimate conditional distributions, we define the causal indicator $\mathcal{C}\in[-1,1]$ [measure_of_causal_strength_oxford, Eells_1991] between an event $X_{i}$ and a label $Y_{j}$ under context $\boldsymbol{Z}$ that we assume fixed for every measurement [measure_of_causal_strength_oxford] as:
$\mathcal{C}(Y_{j},X_{i};\boldsymbol{Z}):=\mathbb{E}_{p(z)}[P(Y_{j}=1\mid X_{i}=1,\boldsymbol{Z}=z)-P(Y_{j}=1\mid X_{i}=0,\boldsymbol{Z}=z)]$ (9)
This enables an easy interpretation, for instance, if $\mathcal{C}<0$, then $X_{i}$ inhibits the occurrence of $Y_{j}$. We employ the term causal indicator to separate from causal strength measures, which, if using this formulation, can be problematic [quantifyingcausalinfluence]. Ours serves more as an indication of the rise in label likelihood after observing a certain event, rather than a strength. The causal strength, strictly speaking, is here the CMI since it serves as a CI-test to recover the correct DAG.
@@ 4 Empirical Evaluation
Comparisons. Although no existing method directly targets one-shot multi-label causal discovery [cd_temporaldata_review], we benchmark OSCAR against local structure learning (LSL) algorithms that estimate global Markov Boundaries. This includes established approaches such as CMB [cmb], MB-by-MB [WANG2014252], PCD-by-PCD [pcdpcd], IAMB [iamb] from the PyCausalFS package [causality_based_feature_selection_2019], as well as the more recent, state-of-the-art MI-MCF [mimcf]. We used a $g4dn.12xlarge$ instance from AWS Sagemaker to run comparisons, containing 4 T4 GPUs. We used a combination of F1-Score, Precision, and Recall with different averaging [reviewmultilabellearning] (Appendix  F.1) to perform the comparisons. The code for OSCAR, $\text{Tf}_{x},\text{Tf}_{y}$ and the evaluation are provided anonymously  [^1] https://github.com/Mathugo/OSCAR-One-Shot-Causal-AutoRegressive-discovery.git as well as the anonymised version of the dataset for reproducibility purposes. An Ablation of the NADEs quality is given in G.1.
Vehicle Event Sequences Dataset. We evaluated our method on a real-world vehicular test set of $n=300,000$ sequences. It contains $|\mathbb{Y}|=474$ different error patterns and about $|\mathbb{X}|=29,100$ different DTCs forming sequences of $\approx 150\pm 90$ events. We used 105m backbones as $\text{Tf}_{x},\text{Tf}_{y}$ [math2024harnessingeventsensorydata]. The two NADEs were not exposed to the test set during. The error patterns are manually defined by domain experts as boolean rules between DTCs in Eq (10) where $(y_{1})$ is a boolean definition based on diagnosis trouble codes $(x_{i})$:
@@
Table 1 shows comparision with $n=50,000$. We found out LSL algorithms failed to compute the Markov Boundaries within multiple days (3-day timeout), far exceeding practical limits for deployment. OSCAR, on the other hand, shows robust classification over a large amount of events $(29,100)$, especially $55\%$ precision, in a matter of minutes. This behaviour highlights the current infeasibility of multi-label causal discovery in high-dimensional event sequences, since it relies on expensive global CI-testings [cd_temporaldata_review]. This positions OSCAR as a more feasible approach for large-scale causal per-sequence causal reasoning in production environments.
@@
Markov Boundary Length. Figure 2 reveals the classification performance depending on the number of nodes in the ground truth MB. On the same plot is drawn in grey the number of samples that each MB length contains (to account for imbalance). We observe that generally, a bigger $|\textbf{MB}(Y_{j})|$ does not imply a reduction in performance, highlighting the capability of OSCAR to retrieve complex Markov Boundaries in high-dimensional data. However, we observe that past a certain number of samples (imbalance threshold in red $\approx 7\times 10^{2}$ samples), the classification metrics are directly correlated with the number of samples per $|\textbf{MB}(Y_{j})|$. This indicates that $\text{Tf}_{y}$ struggles to output proper conditional probabilities, which deteriorates the CI-test when having rare classes. Therefore, when using OSCAR and more generally assumption A4, one should carefully assess class imbalance in the pretraining phase.
Figure 2: Evolution of the One-Shot Recall, Precision and F1-Score in function of the Markov Boundary length $|\textbf{MB}(Y_{j})|$ using $n=45969$ samples.